OpenStudy (anonymous):

the force y measured in newton (N), required to stretch a particular spring x meters as given by y =100x make a table of values using x= 0, 0.5, 1.0 and 1.5 and write the information as ordered pairs

OpenStudy (unklerhaukus):

\[y(x)=100(x)\qquad\qquad(x,y)\] \[y(0)=100(0)=0\qquad\qquad(0,0)\]\[y(0.5)=100(0.5)=50\qquad\qquad(0.5,50)\]\[y(1)=100(1)=\]\[y(1.5)=\]

OpenStudy (anonymous):

Thank you!! How do I explain the meaning of each ordered pair in the context of the problem?

OpenStudy (unklerhaukus):

each orded pair is a solution , the force y is the dependent variable the position is the independent variable